OpenStudy (anonymous):

if f(x)= cosxsin3x, then f'(pi/6) =

OpenStudy (solomonzelman):

can you take the derivative of the expression?

OpenStudy (solomonzelman):

\(\large\color{blue}{ f(x)=\cos(x) \sin^3x }\) \(\large\color{blue}{ f(x)=\cos(x) (\sin x )^3 }\)

OpenStudy (solomonzelman):

Use the product rule, and for the (sin x)^3 do not forget to use the chain.

OpenStudy (solomonzelman):

\(\large\color{blue}{ \frac{\LARGE{dy}}{\LARGE{dx}}=\frac{\LARGE{d}}{\LARGE{dx}}\cos(x) (\sin x )^3+\cos(x)\frac{\LARGE{d}}{\LARGE{dx}} (\sin x)^3 }\)

OpenStudy (solomonzelman):

\(\large\color{blue}{ \frac{\LARGE{dy}}{\LARGE{dx}}=(-\sin~x)(\sin x )^3+\cos(x)\frac{\LARGE{d}}{\LARGE{dx}} (\sin x)^3 }\) \(\large\color{blue}{ \frac{\LARGE{dy}}{\LARGE{dx}}=(-\sin~x)(\sin x )^3+\cos(x)3(\sin x)^2\times \frac{\LARGE{d}}{\LARGE{dx}}\sin(x) }\) \(\large\color{blue}{ \frac{\LARGE{dy}}{\LARGE{dx}}=(-\sin~x)(\sin x )^3+\cos(x)3(\sin x)^2\times(\cos~x) }\)

OpenStudy (solomonzelman):

\(\large\color{blue}{ \frac{\LARGE{dy}}{\LARGE{dx}}=-\sin^4x+3\cos^2(x)\sin ^2x }\)

OpenStudy (anonymous):

is it f'(x)=cosx(sin3x)(cos3x)(3)+sin3x(-sinx)

OpenStudy (solomonzelman):

\(\large\color{blue}{ \frac{\LARGE{dy}}{\LARGE{dx}}=-\sin^4x+3(1-\sin^2x)\sin ^2x }\)

OpenStudy (solomonzelman):

\(\large\color{blue}{ \frac{\LARGE{dy}}{\LARGE{dx}}=-\sin^4x+3(\sin ^2x-\sin^4x)\ }\) \(\large\color{blue}{ \frac{\LARGE{dy}}{\LARGE{dx}}=-\sin^4x+3\sin ^2x-3\sin^4x\ }\) \(\large\color{blue}{ \frac{\LARGE{dy}}{\LARGE{dx}}=-4\sin^4x+3\sin ^2x\ }\)

OpenStudy (solomonzelman):

this is the most simplified derivative expression, now plug in pi/6 for x

OpenStudy (solomonzelman):

sin(30) is 1/2, and you know this from a 30-60-90 triangle. So, \(\large\color{blue}{ f'(pi/6)=-4(1/2)^4+3(1/2)^2\ }\)

OpenStudy (solomonzelman):

all you need is to simplify this expression

OpenStudy (anonymous):

i dont understand how you got dy/dx = -4sin^4x + 3sin^2x

OpenStudy (solomonzelman):

I would suggest to go over my comments, because I was quite clear...

OpenStudy (anonymous):

ok thanks